# MEDIC: Zero-shot Music Editing with Disentangled Inversion Control

### v3

@@ B.3 Data Format
"blended_word": "(\"guitar\", \"violin\")",
"emphasize": "(\"violin\")",
@@
“Editing_prompt" refers to the edited caption, while “source_prompt" denotes the original caption. "Blended_word" indicates the subject to be edited, and “Emphasize" represents the word that should be highlighted. “Editing_instruction" provides a description of the editing process. Additionally, in the editing type “delete instrument," we introduce another key, “neg_prompt", which helps reduce the likelihood of deleted instruments reappearing.
Appendix C Implementation Details
@@
For our evaluation, we utilize the pre-trained AudioLDM 2-Music model (Liu et al., 2023c). Our assessment employs a comprehensive set of metrics, namely CLAP, LPAPS, Structure Distance, and FAD. These metrics are calculated using the CLAP models available in the AudioLDM_eval package, which is accessible at https://github.com/haoheliu/audioldm_eval. In line with the methodology described by  Manor & Michaeli (2024), we apply a forward guidance of 3 and a reverse guidance scale of 12 for DDPM inversion. For the DDIM inversion, the guidance scale is set to 5, while for SDEdit, we employ a guidance scale of 12. The forward guidance of MEDIC is 1 while the reverse scale is 5. We chose these values by exploring different guidance scales, as discussed in Appendix D. We conduct all experiments in NVIDIA 4090. For our evaluation, we have selected the public pre-trained AudioLDM 2-Music model (Liu et al., 2023c). To ensure a thorough and multidimensional assessment, we measure performance using a suite of metrics that includes CLAP, LPAPS, Structure Distance, and FAD and conduct on an NVIDIA 4090. The computation of these metrics is facilitated by the CLAP models provided within the AudioLDM_eval package, which is publicly available at https://github.com/haoheliu/audioldm_eval.
@@
Our methodology is aligned with the protocol established by  Manor & Michaeli (2024), where we have adopted a forward guidance scale of 3 and a reverse guidance scale of 12 for DDPM inversion. In contrast, the DDIM inversion employs a guidance scale of 5, and SDEdit utilizes a guidance scale of 12. For Disentangled Inversion Control, we have determined the forward guidance to be 1 and the reverse scale to be 5. These specific guidance scale values are selected after extensive experimental exploration, the details of which are discussed in Appendix D.
C.1 Metrics
@@
Objective Metrics There are details about four metrics to evaluate the performance of our novel Disentangled Inversion Control framework: (1) CLAP Score (Elizalde et al., 2023): This criterion evaluates the degree to which the output conforms to the specified target prompt. (2) Struture Distance (Ju et al., 2023): Leveraging self-similarity of audio features to measure the structure distance between the source and edited audio. (3) LPAPS (Iashin & Rahtu, 2021; Paissan et al., 2023): An audio adaptation of the Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), this measure evaluates the consistency of the edited audio with the source audio. (4) FAD (Fréchet Audio Distance) (Kilgour et al., 2018): Analogous to the FID used in image analysis, this metric calculates the distance between two distributions of audio signals.
@@ Appendix D Quantitative Results
Table 6: Ablation Studies on Different Guidance Scale
Inference Time We compare the inference time of our method with baselines, and the results are compiled in Table 7. MEDIC achieves the comparative inference time with generation models and inversion techniques. We will make an attempt to reduce the inference time of zero-shot music editing in our future work.
@@
MEDIC 92.0s
Table 7: Inference Time across different methods.
Appendix E Potential Negative Societal Impacts
MEDIC may also lead to potential negative societal impacts that are worthy of consideration. If the data sample of the training model is not diverse enough or biased, the AI-generated music may be overly biased toward one style or element, limiting the diversity of the music and causing discrimination.MEDIC could be used to create fake audio content, such as faking someone’s voice or creating fake musical compositions, posing the risk of fraud and impersonation. Hopefully, all these issues could be taken into consideration when taking the model for real use to avoid ethical issues.
Appendix F Limitations
@@
Appendix G Qualitative Results
For each type in ZoME-Bench, We provide samples to observe the capability of MEDIC intuitively.
G.1 Change Instrument
In Figure 5, we show the capability of MEDIC to change the instrument. Here we edit the ground truth music piece with the source prompt “a live recording of ambient acoustic [guitar] music" and editing prompt “a live recording of ambient acoustic [violin] music". The difference in instruments can be observed in the Mel-spectrum.
@@ G.1 Change Instrument
Figure 5: Editing Type 0 :Change Instrument
G.2 Add Instrument
In Figure 6, we show the capability of MEDIC to add more instruments. Here we edit the ground truth music piece with the source prompt “a heavy metal instructional audio with a distortion guitar" and editing prompt “a heavy metal instructional audio with a distortion guitar [and drums]". The appearance of the new instrument can be observed in the Mel-spectrum which presents a drum sound of high frequency.
@@ G.2 Add Instrument
Figure 6: Editing Type 1 Add Instrument
G.3 Delete Instrument
In Figure 7, we show the capability of MEDIC to delete instruments. Here we edit the ground truth music piece with the source prompt “a lively ska instrumental featuring keyboard trumpets bass [and percussion] with a groovy mood" and the editing prompt “a lively ska instrumental featuring keyboard trumpets and bass with a groovy mood". The vanishing of the instrument can be observed in the Mel-spectrum.
@@ G.3 Delete Instrument
Figure 7: Editing Type 2 Delete Instrument
G.4 Change Genre
In Figure 8, we show the capability of MEDIC to change the genre of a music piece. Here we edit the ground truth music piece with the source prompt “a recording of a solo electric guitar playing [blues] licks" and the editing prompt “a recording of a solo electric guitar playing [rock] licks". The obvious difference in genre can be observed in the Mel-spectrum.
@@ G.4 Change Genre
Figure 8: Editing Type 3 Change Genre
G.5 Change Mood
Mood is an important attribute of music. In Figure 9, we show the capability of MEDIC to change the mood of a music piece. Here we edit the ground truth music piece with the source prompt “a recording of [aggressive] electronic and video game music with synthesizer arrangements" and editing prompt “a recording of [peaceful] electronic and video game music with synthesizer arrangements". The change of mood can be observed in the Mel-spectrum.
@@ G.5 Change Mood
Figure 9: Editing Type 4 Change Mood
G.6 Change Rhythm
Rhythm represents the speed of the music. In Figure 10, we show the capability of MEDIC to change the Rhythm of a music piece. Here we edit the ground truth music piece with the source prompt “a [slow] tempo ukelele tuning recording with static" and the editing prompt “a [fast] tempo ukelele tuning recording with static". The change of Rhythm can be observed in the Mel-spectrum. The edited Mel-spectrum is much more intensive.
@@ G.6 Change Rhythm
Figure 10: Editing Type 5 Change Rhythm
G.7 Change Background
In Figure 11, we show the capability of MEDIC to change the background of the instrument of a music piece. Here we edit the ground truth music piece with the source prompt “an amateur ukulele recording with a [medium to uptempo] pace" and editing prompt “an amateur ukulele recording with a [steady and rhythmic] pace". The change of instrument background can be observed in the Mel-spectrum.
source: an amateur ukulele recording with a [medium to uptempo] pace" and editing prompt
@@ G.7 Change Background
Figure 11: Editing Type 6 Change Background
G.8 Change Melody
In Figure 12, we show the capability of MEDIC to change the melody of a music piece. Here we edit the ground truth music piece with the source prompt “this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo" and editing prompt “this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo". The change of Melody can be observed in the Mel-spectrum.
@@ G.8 Change Melody
Figure 12: Editing Type 7 Change Melody
G.9 Extract Instrument
In Figure 13, we show the capability of MEDIC to extract one certain instrument of a music piece. Here we edit the ground truth music piece with the source prompt “a reggae rhythm recording with bongos [djembe congas acoustic drums and electric guitar]" and editing prompt “a reggae rhythm recording with bongos". The change of instruments can be observed in the Mel-spectrum.
@@ G.9 Extract Instrument
Figure 13: Editing Type 8 Extract Instrument
Appendix H safeguards

### v4

@@ 7.3. Data Format
"blended_word": "("guitar", "violin")",
"emphasize": "("violin")",
@@
“Editing_prompt” refers to the edited caption, while “source_prompt” denotes the original caption. ”Blended_word” indicates the subject to be edited, and “Emphasize” represents the word that should be highlighted. “Editing_instruction” provides a description of the editing process. Additionally, in the editing type “delete instrument,” we introduce another key, “neg_prompt”, which helps reduce the likelihood of deleted instruments reappearing.
8. Implementation Details
@@
For our evaluation, we utilize the pre-trained AudioLDM 2-Music model (liu2023audioldm2). Our assessment employs a comprehensive set of metrics, namely CLAP, LPAPS, Structure Distance, and FAD. These metrics are calculated using the CLAP models available in the AudioLDM_eval package, which is accessible at https://github.com/haoheliu/audioldm_eval. In line with the methodology described by  manor2024zero, we apply a forward guidance of 3 and a reverse guidance scale of 12 for DDPM inversion. For the DDIM inversion, the guidance scale is set to 5, while for SDEdit, we employ a guidance scale of 12. The forward guidance of MelodyEdit is 1 while the reverse scale is 5. We chose these values by exploring different guidance scales, as discussed in Appendix 9. We conduct all experiments in NVIDIA 4090.
@@
Our methodology is aligned with the protocol established by  manor2024zero, where we have adopted a forward guidance scale of 3 and a reverse guidance scale of 12 for DDPM inversion. In contrast, the DDIM inversion employs a guidance scale of 5, and SDEdit utilizes a guidance scale of 12. For Disentangled Inversion Control, we have determined the forward guidance to be 1 and the reverse scale to be 5. These specific guidance scale values are selected after extensive experimental exploration, the details of which are discussed in Appendix 9.
8.1. Metrics
@@
Objective Metrics There are details about four metrics to evaluate the performance of our novel Disentangled Inversion Control framework: (1) CLAP Score (elizalde2023clap): This criterion evaluates the degree to which the output conforms to the specified target prompt. (2) Struture Distance (ju2023direct): Leveraging self-similarity of audio features to measure the structure distance between the source and edited audio. (3) LPAPS (iashin2021taming; paissan2023audio): An audio adaptation of the Learned Perceptual Image Patch Similarity (LPIPS) (zhang2018unreasonable), this measure evaluates the consistency of the edited audio with the source audio. (4) FAD (Fréchet Audio Distance) (kilgour2018fad): Analogous to the FID used in image analysis, this metric calculates the distance between two distributions of audio signals. (5) Chroma (Chromagram Similarity) (zhang2024musicmagus): The average cosine similarity between the chromagrams of the original and edited music, which denotes the preservation of pitch contours and rhythm patterns in the music. (6) Accuracy (chu2024qwen2): The rate of successful editing judged by the Qwen model, calculated by constructing a question-answer (QA) pair, where the model’s output is compared against the expected result. The comparison assesses whether the model has made the correct edit.
@@ 9. Quantitative Results
Table 5. Ablation Studies on Different Guidance Scale
Inference Time We compare the inference time of our method with baselines, and the results are compiled in Table 6. MelodyEdit achieves the comparative inference time with generation models and inversion techniques. We will make an attempt to reduce the inference time of zero-shot music editing in our future work.
@@
MelodyEdit 92.0s
Table 6. Inference Time across different methods.
10. Potential Negative Societal Impacts
MelodyEdit may also lead to potential negative societal impacts that are worthy of consideration. If the data sample of the training model is not diverse enough or biased, the AI-generated music may be overly biased toward one style or element, limiting the diversity of the music and causing discrimination.MelodyEdit could be used to create fake audio content, such as faking someone’s voice or creating fake musical compositions, posing the risk of fraud and impersonation. Hopefully, all these issues could be taken into consideration when taking the model for real use to avoid ethical issues.
11. Limitations
@@
12. Qualitative Results
For each type in ZoME-Bench, We provide samples to observe the capability of MelodyEdit intuitively.
12.1. Change Instrument
In Figure 7, we show the capability of MelodyEdit to change the instrument. Here we edit the ground truth music piece with the source prompt “a live recording of ambient acoustic [guitar] music” and editing prompt “a live recording of ambient acoustic [violin] music”. The difference in instruments can be observed in the Mel-spectrum.
@@ 12.1. Change Instrument
Figure 7. Editing Type 0 :Change Instrument
12.2. Add Instrument
In Figure 8, we show the capability of MelodyEdit to add more instruments. Here we edit the ground truth music piece with the source prompt “a heavy metal instructional audio with a distortion guitar” and editing prompt “a heavy metal instructional audio with a distortion guitar [and drums]”. The appearance of the new instrument can be observed in the Mel-spectrum which presents a drum sound of high frequency.
@@ 12.2. Add Instrument
Figure 8. Editing Type 1 Add Instrument
12.3. Delete Instrument
In Figure 9, we show the capability of MelodyEdit to delete instruments. Here we edit the ground truth music piece with the source prompt “a lively ska instrumental featuring keyboard trumpets bass [and percussion] with a groovy mood” and the editing prompt “a lively ska instrumental featuring keyboard trumpets and bass with a groovy mood”. The vanishing of the instrument can be observed in the Mel-spectrum.
@@ 12.3. Delete Instrument
Figure 9. Editing Type 2 Delete Instrument
12.4. Change Genre
In Figure 10, we show the capability of MelodyEdit to change the genre of a music piece. Here we edit the ground truth music piece with the source prompt “a recording of a solo electric guitar playing [blues] licks” and the editing prompt “a recording of a solo electric guitar playing [rock] licks”. The obvious difference in genre can be observed in the Mel-spectrum.
@@ 12.4. Change Genre
Figure 10. Editing Type 3 Change Genre
12.5. Change Mood
Mood is an important attribute of music. In Figure 11, we show the capability of MelodyEdit to change the mood of a music piece. Here we edit the ground truth music piece with the source prompt “a recording of [aggressive] electronic and video game music with synthesizer arrangements” and editing prompt “a recording of [peaceful] electronic and video game music with synthesizer arrangements”. The change of mood can be observed in the Mel-spectrum.
@@ 12.5. Change Mood
Figure 11. Editing Type 4 Change Mood
12.6. Change Rhythm
Rhythm represents the speed of the music. In Figure 12, we show the capability of MelodyEdit to change the Rhythm of a music piece. Here we edit the ground truth music piece with the source prompt “a [slow] tempo ukelele tuning recording with static” and the editing prompt “a [fast] tempo ukelele tuning recording with static”. The change of Rhythm can be observed in the Mel-spectrum. The edited Mel-spectrum is much more intensive.
@@ 12.6. Change Rhythm
Figure 12. Editing Type 5 Change Rhythm
12.7. Change Background
In Figure 13, we show the capability of MelodyEdit to change the background of the instrument of a music piece. Here we edit the ground truth music piece with the source prompt “an amateur ukulele recording with a [medium to uptempo] pace” and editing prompt “an amateur ukulele recording with a [steady and rhythmic] pace”. The change of instrument background can be observed in the Mel-spectrum.
source: an amateur ukulele recording with a [medium to uptempo] pace” and editing prompt
@@ 12.7. Change Background
Figure 13. Editing Type 6 Change Background
12.8. Change Melody
In Figure 14, we show the capability of MelodyEdit to change the melody of a music piece. Here we edit the ground truth music piece with the source prompt “this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo” and editing prompt “this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo”. The change of Melody can be observed in the Mel-spectrum.
@@ 12.8. Change Melody
Figure 14. Editing Type 7 Change Melody
12.9. Extract Instrument
In Figure 15, we show the capability of MelodyEdit to extract one certain instrument of a music piece. Here we edit the ground truth music piece with the source prompt “a reggae rhythm recording with bongos [djembe congas acoustic drums and electric guitar]” and editing prompt “a reggae rhythm recording with bongos”. The change of instruments can be observed in the Mel-spectrum.
@@ 12.9. Extract Instrument
Figure 15. Editing Type 8 Extract Instrument
13. safeguards